# Non-Hermitian symmetry breaking and Lee-Yang theory for quantum XYZ and clock models

### v1

@@ Abstract
Lee-Yang theory offers a unifying framework for understanding classical phase transitions and dynamical quantum phase transitions through the analysis of partition functions and Loschmidt echoes. Recently, this framework is extended to characterize quantum phase transitions in arXiv:2509.20258 by introducing the concepts of non-Hermitian symmetry breaking and fidelity zeros. Here, we generalize the theory by studying a broad class of quantum models, including the XY model, the XXZ model, the XYZ model, and the $\mathbb{Z}_{3}$ clock model in one dimension, subject to complex external magnetic field. For the XY, XXZ and XYZ models, we find that the complex field breaks parity symmetry and induces oscillations of the ground state between the two parity sectors, giving rise to fidelity zeros within the ordered phases. For the $\mathbb{Z}_{3}$ clock model, the complex field splits the real part of the ground-state energy between the neutral sector ($q=0$) and the charged sectors ($q=1,2$), while preserving the degeneracy within the charged sector. Fidelity zeros arise only after projecting out one of the charged sectors, and the finite-size scaling of these zeros produces critical exponents fully consistent with analytical predictions.
@@ I Introduction
Understanding phase transitions is one of the central topics in physics [sachdev1999quantum]. The logarithm of the partition function, which is equivalent to the Helmholtz free energy in the canonical ensemble, serves as the central thermodynamic potential, with its nonanalytic behavior marking the emergence of a phase transition [bena2005statistical, wang2025extracting]. However, because the partition function is composed of statistical weights that are analytic functions, its logarithm cannot develop any nonanalyticity in a finite system [bena2005statistical]. To understand how phase transitions actually occur, Lee and Yang studied the partition function in finite systems under a complex magnetic field and developed a framework [yang1952statistical, lee1952statistical] for characterizing phase transitions by analyzing its zeros, now known as Lee-Yang zeros [lee1952statistical]. For the ferromagnetic Ising model, Lee and Yang proved that all zeros lie precisely on the unit circle [lee1952statistical, bena2005statistical, wei2012lee].
@@
This compelling analysis of phase transitions and the distribution of zeros has attracted both theoretical [bena2005statistical, heyl2018dynamical] and experimental [peng2015experimental, brandner2017experimental] interest over the past decades. The beauty of the Lee-Yang theory lies in identifying the zeros of the partition function in finite systems under a complex magnetic field, which appear only in systems with a real field in the thermodynamic limit. Fisher later extended this idea to study phase transitions by analyzing the zeros of the partition function (Fisher zeros) in the complex temperature plane [fisher1965statistical]. Indeed, the key idea of analyzing the zeros of the partition function under a complex control parameter is quite general. For example, beyond the Lee-Yang zeros and Fisher zeros, it has been shown that phase transitions can be explored through complex $q$ in the $q$-state Potts model, giving rise to the concept of Potts zeros [kim2001density]. Moreover, for quantum systems subjected to a sudden quench, the Loschmidt echo, which is analogous to the partition function upon defining the imaginary time $\tau=it$ [heyl2013dynamical], is introduced to probe the system’s dynamical behavior. The nonequilibrium phenomenon known as the dynamical quantum phase transition [heyl2013dynamical] arises in the time plane at the zeros of the Loschmidt echo.
@@
The distribution of zeros is another research interest. Originally, the Lee-Yang unit circle theorem was discovered in the ferromagnetic Ising model [lee1952statistical]. It was subsequently extended to general Heisenberg spin models, showing that the theorem holds as long as the $z$-$z$ coupling is ferromagnetic and dominant [suzuki1969distribution, asano1970lee, asano1970theorems, suzuki1971zeros]. The distribution of zeros on the unit circle implies that they concentrate along the imaginary axis of the magnetic field. Furthermore, it is shown that at high temperatures $T>T_{c}$, a gap develops in the distribution, giving rise to branch points, referred to as the Yang-Lee edge singularity [fisher1978yang, kurtze1979yang, cardy1985conformal, cardy1989s, wei2017probing]. With decreasing temperature, the zeros approach and pinch the real axis [ananikian2015imaginary], signaling a phase transition at $T=T_{c}$ in the thermodynamic limit. Recently, Lee-Yang zeros and the Yang-Lee edge singularity have been revisited both theoretically [yin2017kibble, jian2021yang, sanno2022engineering, shen2023proposal, lu2025dynamical, xu2025characterizing, zhang2025yang, sun2025hybrid] and experimentally [gao2024experimental, lan2024experimental] because of recent development in non-Hermitian physics.
@@
In addition to classical and dynamical quantum phase transitions, quantum phase transitions driven by quantum fluctuations are a central topic of broad interest in condensed matter physics [sachdev1999quantum, sondhi1997continuous, vojta2003quantum]. A natural extension of the Lee-Yang framework to quantum systems is to compute the partition function (or generating functions) of quantum many-body systems and analyze the distribution of zeros [kist2021lee, vecsei2022lee, vecsei2023lee, liu2024imaginary, liu2024exact, vecsei2025lee, li2023yang, li2025yang, meng2025detecting, wang2024quantum, he2025revisiting]. Nevertheless, it is well-known that the most direct approach to quantum phase transitions is through the ground state and its associated observables. A key question is whether the Lee-Yang framework can be captured from the ground states of quantum many-body systems. This question has been addressed in recent work [gu2025fidelity] based on the concepts of non-Hermitian symmetry breaking and fidelity (overlap of two ground states) zeros. These zeros, examined in one- and two-dimensional ferromagnetic Ising models [gu2025fidelity], confirm the validity of the Lee-Yang theory for quantum phase transitions via the ground states in the complex magnetic field plane. Another key question is whether this fidelity-based Lee-Yang framework applies solely to the ferromagnetic Ising model or can be extended to a broad class of models, as in the original Lee-Yang theory for classical phase transitions.
@@
In this work, we extend the framework introduced in Ref. [gu2025fidelity] to general quantum XYZ and $\mathbb{Z}_{3}$ clock models. For the XYZ model, we show that the Lee-Yang theory holds by examining the XY and XXZ models, where parity is broken by the complex field, as in the Ising model. In the $\mathbb{Z}_{3}$ clock model, the $\mathbb{Z}_{3}$ symmetry is similarly broken under the complex field. However, since the complex field cannot lift the degeneracy of ground states in the charged sectors ($q=1,2$), the fidelity-based Lee-Yang theory can only be formulated by projecting out one of these sectors. This work therefore generalizes the fidelity-based Lee-Yang theory of Ref. [gu2025fidelity] to broader classes of models and higher discrete symmetries, demonstrating the wide applicability of the Lee-Yang framework.
@@ II XYZ model
The general anisotropic Heisenberg Hamiltonian for the XYZ model with transverse and longitudinal magnetic fields is given by [asano1970lee, asano1970theorems, suzuki1971zeros],
@@ II.1 XY model
We first consider the ferromagnetic XY model under a complex transverse field applied along the $z$ direction by setting $J_{x},J_{y}>0$, $J_{z}=0$ and $h_{x}=0$. The Hamiltonian of the anisotropic XY model is given by [pi2021phase],
@@
The non-Hermitian XY model can be solved exactly by mapping spin operators onto spinless fermions [pi2021phase]. To this end, it is convenient to rewrite Hamiltonian (2) using the spin ladder operators $\sigma_{i}^{\pm}=\frac{1}{2}(\sigma_{i}^{x}\pm i\sigma_{i}^{y})$. The Hamiltonian (2) then becomes,
@@
The fidelity between $h$ and $h^{\prime}$ is given by [gu2025fidelity],
@@
We analytically calculate the ground-state fidelity of the XY model by varying the real part of the complex magnetic field $h$ for $L=16$ and $\gamma=0.8$, and show the corresponding distribution of fidelity zeros in Fig.1, with $J=1$ fixed throughout the numerical simulations. As illustrated in Fig. 1(a), fidelity zeros appear in the system when $\text{Re}(h)<\text{Re}(h_{L})$, while no zeros are observed for $\text{Re}(h)>\text{Re}(h_{L})$. This behavior indicates that a quantum phase transition occurs at $\text{Re}(h_{L})$. To study the finite-size scaling of fidelity zeros near the critical point, we first define $h_{L}$ as the complex field $h$ with the largest real part and the smallest imaginary part in our finite-size analysis. The finite-size scaling of both $\text{Re}(h_{L})$ and $\text{Im}(h_{L})$ is analyzed using the relation [gu2025fidelity],
@@
We further examine the distribution of fidelity zeros in the complex magnetic field $h=ge^{i\theta}$ for $L=16$ [c.f. Fig.1(c) and (d)]. These zeros are computed for $g=0.5$ and $g=1.5$, while $\theta$ is varied over the interval $(0,2\pi]$. By introducing the fugacity $Z=e^{i\theta}$, we observe that the zeros lie on the unit circle in the complex-$h$ plane. For $g<h_{c}$, they are uniformly distributed along the circle; in contrast, for $g>h_{c}$, they condense into two circular arcs separated by a well-defined gap. These observations confirm that the Lee-Yang theory also holds for the XY model, as previously demonstrated for the transverse field Ising model [gu2025fidelity].
@@
In the case $J_{x}=J_{y}\neq J_{z}$ and $h_{z}=0$, the system reduces to the XXZ model under a transverse field along the $x$ direction. When $J_{x}=J_{y}<0$ and $J_{z}/J_{x}<-1$, the system undergoes a phase transition from a ferromagnetic to a paramagnetic phase [dmitriev2002gap]. Furthermore, for the general XYZ model with $J_{x}\neq J_{y}\neq J_{z}$ and $h_{z}=0$, a ferromagnetic-to-paramagnetic phase transition also occurs when the coupling $J_{z}>0$ is dominant. Numerical simulations of the XXZ and XYZ models, presented in Fig.2, further support the validity of the Lee-Yang framework in characterizing quantum phase transitions in general spin systems.
@@ III Clock model
We have shown that the fidelity-based Lee-Yang framework is valid for the general spin-1/2 XYZ model. In the following, we extend this theory to the $\mathbb{Z}_{3}$ clock model. The one-dimensional quantum $\mathbb{Z}_{3}$ clock model is described by the Hamiltonian [ortiz2012dualities, chen2017phase, sun2019phase, tang2023dynamical, yu2023dynamical, yu2024emergent]:
@@
where $J$ denotes the coupling strength and and $h$ represents the complex transverse field. The operators $V_{j}$ and $U_{j}$ are defined by,
@@
respectively, with $\omega=e^{2\pi i/3}$ and $(V_{j})^{3}=(U_{j})^{3}=\mathbb{I}$. Here, we impose $J=1$ and periodic boundary conditions $V_{L+1}=V_{1}$. For a real field $\text{Im(h)}=0$, the $\mathbb{Z}_{3}$ clock model undergoes a phase transition at the critical point $\text{Re}(h)=1$, from a $\mathbb{Z}_{3}$-symmetric gapped ordered phase to a gapped disordered phase, with the correlation-length critical exponent $\nu=5/6$ [sun2019phase]. In the ordered phase, the three degenerate ground states lie in distinct symmetry sectors, determined by the many-body $\mathbb{Z}_{3}$ projector,
$P_{q}=\frac{1}{3}\left[\mathbb{I}+\omega^{-q}\prod_{j}U_{j}+\omega^{-2q}\left(\prod_{j}U_{j}\right)^{2}\right],$ (16)
@@
However, when the external field is complex, for example $h=ge^{i\theta}$, it leads to the non-Hermitian symmetry-breaking and lifts the ground-state degeneracy, defined by the lowest real part of the complex energies. Notably, the complex field only lifts the degeneracy between the neutral sector ($q=0$) and the charged sectors ($q=1,2$), while the two ground-state energies within the charged sectors ($q=1,2$) remain degenerate, as shown in Fig.3. Because of this degeneracy, the ground state cannot be uniquely assigned to either the $q=1$ or $q=2$ sector. Consequently, fidelity is ill-defined, as the ground state forms a superposition of states from the $q=1$ or $q=2$ sectors. To lift this degeneracy and select a unique charged sector, we introduce an additional projector $P^{\prime}=\mathbb{I}-P_{2}$, which completely projects out the $q=2$ sector. This results in a clear non-Hermitian symmetry-breaking between the $q=0$ and $q=1$ sector, which can be detected through fidelity zeros.
Figure 4: Fidelity zeros and fidelity edges in the clock model. (a) Distribution of fidelity zeros in the complex-$h$ plane for $L=10$. The scatter points mark the complex field values at which the fidelity vanishes, with the critical point located at $Re(h)\approx 1$. (b) Finite-size scaling of the fidelity zeros as a function of system size $L$ from 6 to 15. The green dots denote the complex field values with the largest real part and smallest imaginary part of the fidelity zeros; the black dashed line shows the fitted curve, and the red star on the real axis marks the critical value $h_{c}=1.006$. (c)-(d) Fidelity-zero distributions in the complex-field plane for $L=10$ for $g=0.5$ and $g=2.5$, respectively. All zeros lie on the unit circle $Z=e^{i\theta}$ with $\theta=(0,2\pi]$; Fidelity edges appear for $g=2.5$.
@@
To validate our results, we numerically compute the fidelity zeros for a system of size $L=10$ in the complex field plane. The zeros are found exclusively within the ordered phase and vanish abruptly in the disordered phase, as shown in Fig.4(a). For $L=10$, the phase transition is located at $\text{Re}(h_{L})=1.0238$. As the lattice size increases, the real part of $h$ approaches the real axis while the imaginary part vanishes in the thermodynamic limit, yielding a critical point $h_{c}=1.006$ with $\nu=5/6$, as illustrated in Fig.4(b) through the finite-size scaling according to Eq.(13). Furthermore, under the condition $h=gZ=e^{i\theta}$, all fidelity zeros lie on the unit circle in the complex-$Z$ plane, with a total of $2L$ zeros. Their distribution, however, differs significantly between the ordered and disordered phases. For $g<h_{c}$, the zeros are distributed almost uniformly along the circle [c.f. Fig.4(c)], whereas for $g<h_{c}$, they coalesce into two distinct arcs [c.f. Fig.4(d)]. These results demonstrate that our approach is applicable to models with higher discrete symmetries, such as extending from $\mathbb{Z}_{2}$ to $\mathbb{Z}_{3}$.

### v2

@@ Abstract
Lee-Yang theory offers a unifying framework for understanding classical phase transitions and dynamical quantum phase transitions through the analysis of partition functions and Loschmidt echoes. Recently, this framework is extended to characterize quantum phase transitions of quantum Ising models by introducing the concepts of non-Hermitian parity-symmetry breaking and fidelity zeros. Here, we generalize the theory by studying a broad class of quantum models, including the XY, the XXZ, the XYZ, and the $\mathbb{Z}_{p}$ clock models in one dimension, subject to a complex magnetic field. For the XY, XXZ and XYZ models, we find that the complex field breaks parity symmetry and induces oscillations of the ground state between the two parity sectors, giving rise to fidelity zeros within the ordered phases. For the $\mathbb{Z}_{3}$ clock model, the complex field splits the real part of the ground-state energy between the neutral sector ($q=0$) and the charged sectors ($q=1,2$), while preserving the degeneracy within the charged sector. Fidelity zeros arise only after projecting out one of the charged sectors. In contrast, for the $\mathbb{Z}_{4}$ clock model, the ground state oscillates between the neutral sector ($q=0$) and the charged sector ($q=2$), which directly gives rise to fidelity zeros. Finite-size scaling of these zeros yields critical exponents in full agreement with analytical predictions, demonstrating that this approach is applicable not only to the Ising model with $\mathbb{Z}_{2}$ symmetry, but also to more general Heisenberg-type models and systems with higher discrete symmetries.
@@ I Introduction
Understanding phase transitions is one of the central topics in physics [34]. The logarithm of the partition function, which is equivalent to the Helmholtz free energy in the canonical ensemble, serves as the central thermodynamic potential, with its nonanalytic behavior marking the emergence of a phase transition [4, 49]. However, because the partition function is composed of statistical weights that are analytic functions, its logarithm cannot develop any nonanalyticity in a finite system [4]. To understand how phase transitions actually occur, Lee and Yang studied the partition function in finite systems under a complex magnetic field and developed a framework [54, 22] for characterizing phase transitions by analyzing its zeros, now known as Lee-Yang zeros [22]. For the ferromagnetic Ising model, Lee and Yang proved that all zeros lie precisely on the unit circle [22, 4, 50].
@@
This compelling analysis of phase transitions and the distribution of zeros has attracted both theoretical [4, 16] and experimental [32, 5] interest over the past decades. The beauty of the Lee-Yang theory lies in identifying the zeros of the partition function in finite systems under a complex magnetic field, which appear only in systems with a real field in the thermodynamic limit. Fisher later extended this idea to study phase transitions by analyzing the zeros of the partition function (Fisher zeros) in the complex temperature plane [10]. Indeed, the key idea of analyzing the zeros of the partition function under a complex control parameter is quite general. For example, beyond the Lee-Yang zeros and Fisher zeros, it has been shown that phase transitions can be explored through complex $q$ in the $q$-state Potts model, giving rise to the concept of Potts zeros [18]. Moreover, for quantum systems subjected to a sudden quench, the Loschmidt echo, which is analogous to the partition function upon defining the imaginary time $\tau=it$ [15], is introduced to probe the system’s dynamical behavior. The nonequilibrium phenomenon known as the dynamical quantum phase transition [15] arises in the time plane at the zeros of the Loschmidt echo.
@@
The distribution of zeros is another research interest. Originally, the Lee-Yang unit circle theorem was discovered in the ferromagnetic Ising model [22]. It was subsequently extended to general Heisenberg spin models, showing that the theorem holds as long as the $z$-$z$ coupling is ferromagnetic and dominant [41, 2, 3, 40]. The distribution of zeros on the unit circle implies that they concentrate along the imaginary axis of the magnetic field. Furthermore, it is shown that at high temperatures $T>T_{c}$, a gap develops in the distribution, giving rise to branch points, referred to as the Yang-Lee edge singularity [11, 20, 7, 6, 51]. With decreasing temperature, the zeros approach and pinch the real axis [1], signaling a phase transition at $T=T_{c}$ in the thermodynamic limit. Recently, Lee-Yang zeros and the Yang-Lee edge singularity have been revisited both theoretically [57, 56, 17, 35, 36, 55, 31, 27, 52, 60, 39, 53] and experimentally [12, 21].
@@
In addition to classical and dynamical quantum phase transitions, quantum phase transitions [34, 37, 47] driven by quantum fluctuations are a central topic of broad interest in condensed matter physics. A natural extension of the Lee-Yang framework to quantum systems is to compute the partition function (or generating functions) of quantum many-body systems and analyze the distribution of zeros [43, 19, 45, 44, 25, 26, 46, 23, 24, 29, 48, 14, 28]. Nevertheless, it is well-known that the most direct approach to quantum phase transitions is through the ground state and its associated observables. A key question is whether the Lee-Yang framework can be captured from the ground states of quantum many-body systems. This question has been addressed in recent work [13] based on the concepts of non-Hermitian symmetry breaking and fidelity (overlap of two ground states) zeros. These zeros, examined in one- and two-dimensional ferromagnetic Ising models [13], confirm the validity of the Lee-Yang theory for quantum phase transitions via the ground states in the complex magnetic field plane. However, an important open question remains: does this fidelity-based Lee-Yang framework apply only to the ferromagnetic Ising model, or can it be extended to a broader class of quantum many-body systems, analogous to the generality of the original Lee-Yang theory for classical phase transitions?
@@
In this work, we extend the framework developed in Ref. [13] to general quantum XYZ and $\mathbb{Z}_{p}$ clock models. For the XYZ model, we show that the Lee-Yang theory holds by examining the XY, XXZ and XYZ models, where parity is broken by the complex field, as in the Ising model. In the $\mathbb{Z}_{p}$ clock model, the $\mathbb{Z}_{p}$ symmetry is similarly broken under the complex field. However, since the complex field cannot lift the degeneracy of ground states in the charged sectors ($q=1,2$), the fidelity-based Lee-Yang theory can only be formulated by projecting out one of these sectors in the $p=3$ clock model. This work therefore generalizes the fidelity-based Lee-Yang theory to broader classes of models and higher discrete symmetries, demonstrating the wide applicability of the Lee-Yang framework.
@@ II XYZ model
The general anisotropic Heisenberg Hamiltonian for the XYZ model with transverse and longitudinal magnetic fields is given by [2, 3, 40],
@@ II.1 XY model
We first consider the ferromagnetic XY model under a complex transverse field applied along the $z$ direction by setting $J_{x},J_{y}>0$, $J_{z}=0$ and $h_{x}=0$. The Hamiltonian of the anisotropic XY model is given by [33],
@@
The non-Hermitian XY model can be solved exactly by mapping spin operators onto spinless fermions [33]. To this end, it is convenient to rewrite Hamiltonian (2) using the spin ladder operators $\sigma_{i}^{\pm}=\frac{1}{2}(\sigma_{i}^{x}\pm i\sigma_{i}^{y})$. The Hamiltonian (2) then becomes,
@@
The fidelity between $h$ and $h^{\prime}$ is given by [13],
@@
We analytically calculate the ground-state fidelity of the XY model by varying the real part of the complex magnetic field $h$ for $L=16$ and $\gamma=0.8$, and show the corresponding distribution of fidelity zeros in Fig. 1, with $J=1$ fixed throughout the numerical simulations. As illustrated in Fig. 1(a), fidelity zeros appear in the system when $\text{Re}(h)<\text{Re}(h_{L})$, while no zeros are observed for $\text{Re}(h)>\text{Re}(h_{L})$. This behavior indicates that a quantum phase transition occurs at $\text{Re}(h_{L})$. To study the finite-size scaling of fidelity zeros near the critical point, we first define $h_{L}$ as the complex field $h$ with the largest real part and the smallest imaginary part in our finite-size analysis. The finite-size scaling of both $\text{Re}(h_{L})$ and $\text{Im}(h_{L})$ is analyzed using the relation [13],
@@
We further examine the distribution of fidelity zeros in the complex magnetic field $h=ge^{i\theta}$ for $L=16$ [c.f. Fig. 1(c) and (d)]. These zeros are computed for $g=0.5$ and $g=1.5$, while $\theta$ is varied over the interval $(0,2\pi]$. By introducing the fugacity $Z=e^{i\theta}$, we observe that the zeros lie on the unit circle in the complex-$h$ plane. For $g<h_{c}$, they are uniformly distributed along the circle; in contrast, for $g>h_{c}$, they condense into two circular arcs separated by a well-defined gap. These observations confirm that the Lee-Yang theory also holds for the XY model, as previously demonstrated for the transverse field Ising model [13].
@@
In the case $J_{x}=J_{y}\neq J_{z}$ and $h_{z}=0$, the system reduces to the XXZ model under a transverse field along the $x$ direction. When $J_{x}=J_{y}<0$ and $J_{z}/J_{x}<-1$, the system undergoes a phase transition from a ferromagnetic to a paramagnetic phase [9]. Furthermore, for the general XYZ model with $J_{x}\neq J_{y}\neq J_{z}$ and $h_{z}=0$, a ferromagnetic-to-paramagnetic phase transition also occurs when the coupling $J_{z}>0$ is dominant. Numerical simulations of the XXZ and XYZ models, presented in Fig. 2, further support the validity of the Lee-Yang framework in characterizing quantum phase transitions in general spin systems.
@@ III Clock model
We have shown that the fidelity-based Lee-Yang framework is valid for the general spin-1/2 XYZ model. In the following, we extend this theory to the $\mathbb{Z}_{p}$ clock model. The one-dimensional quantum $\mathbb{Z}_{p}$ clock model is described by the Hamiltonian [30, 8, 38, 42, 59, 58]:
@@
where $J$ denotes the coupling strength and and $h$ represents the complex transverse field. The quantum clock model exhibits a second-order phase transition for $p\leq 4$, whereas it undergoes a Berezinskii–Kosterlitz–Thouless transition for $p>5$ [38]. In the following, we restrict our analysis to the second-order phase-transition regime. For $p=2$, the quantum clock model is exactly equivalent to the transverse-field Ising chain. For $p=3$, the operators $V_{j}$ and $U_{j}$ are defined by,
@@
respectively, where $\omega=e^{2\pi i/3}$ and $(V_{j})^{3}=(U_{j})^{3}=\mathbbm{1}$. Here, we impose $J=1$ and periodic boundary conditions $V_{L+1}=V_{1}$. For a real field $\text{Im(h)}=0$, the $\mathbb{Z}_{3}$ clock model undergoes a phase transition at the critical point $\text{Re}(h)=1$, from a $\mathbb{Z}_{3}$-symmetric gapped ordered phase to a gapped disordered phase, with the correlation-length critical exponent $\nu=5/6$ [38].
Figure 4: Fidelity zeros and fidelity edges in the clock model. (a) Distribution of fidelity zeros in the complex-$h$ plane for $L=10$. The scatter points mark the complex field values at which the fidelity vanishes, with the critical point located at $Re(h)\approx 1$. (b) Finite-size scaling of the fidelity zeros as a function of system size $L$ from 7 to 15. The green dots denote the complex field values with the largest real part and smallest imaginary part of the fidelity zeros; the black dashed line shows the fitted curve, and the red star on the real axis marks the critical value $h_{c}=1.006$. (c)-(d) Fidelity-zero distributions in the complex-field plane for $L=10$ for $g=0.5$ and $g=2.5$, respectively. All zeros lie on the unit circle $Z=e^{i\theta}$ with $\theta=(0,2\pi]$; Fidelity edges appear for $g=2.5$.
In the ordered phase, the three degenerate ground states lie in distinct symmetry sectors, determined by the many-body $\mathbb{Z}_{3}$ projector,
$P_{q}=\frac{1}{3}\left[\mathbbm{1}+\omega^{-q}\prod_{j}U_{j}+\omega^{-2q}\left(\prod_{j}U_{j}\right)^{2}\right],$ (16)
@@
However, when the external field is complex, for example $h=ge^{i\theta}$, it leads to the non-Hermitian symmetry-breaking and lifts the ground-state degeneracy, defined by the lowest real part of the complex energies. Notably, the complex field only lifts the degeneracy between the neutral sector ($q=0$) and the charged sectors ($q=1,2$), while the two ground-state energies within the charged sectors ($q=1,2$) remain degenerate, as shown in Fig. 3. Because of this degeneracy, the ground state cannot be uniquely assigned to either the $q=1$ or $q=2$ sector. Consequently, fidelity is ill-defined, as the ground state forms a superposition of states from the $q=1$ or $q=2$ sectors. To lift this degeneracy and select a unique charged sector, we introduce an additional projector $P^{\prime}=\mathbbm{1}-P_{2}$, which completely projects out the $q=2$ sector. This results in a clear non-Hermitian symmetry-breaking between the $q=0$ and $q=1$ sector, which can be detected through fidelity zeros.
@@
To validate our results, we numerically compute the fidelity zeros for a system of size $L=10$ in the complex field plane. The zeros are found exclusively within the ordered phase and vanish abruptly in the disordered phase, as shown in Fig. 4(a). For $L=10$, the phase transition is located at $\text{Re}(h_{L})=1.0238$. As the lattice size increases, the real part of $h$ approaches the real axis while the imaginary part vanishes in the thermodynamic limit, yielding a critical point $h_{c}=1.006$ with $\nu=5/6$, as illustrated in Fig. 4(b) through the finite-size scaling according to Eq. (13). Furthermore, under the condition $h=gZ=e^{i\theta}$, all fidelity zeros lie on the unit circle in the complex-$Z$ plane, with a total of $2L$ zeros. Their distribution, however, differs significantly between the ordered and disordered phases. For $g<h_{c}$, the zeros are distributed almost uniformly along the circle [c.f. Fig. 4(c)], whereas for $g<h_{c}$, they coalesce into two distinct arcs [c.f. Fig. 4(d)]. These results demonstrate that fidleity-based Lee-Yang approach is applicable to models with higher $\mathbb{Z}_{3}$ symmetries.
For $p=4$ quantum clock model, the operators $V_{j}$ and $U_{j}$ are given by
$V_{j}=\begin{pmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\end{pmatrix},U_{j}=\begin{pmatrix}1&0&0&0\\ 0&\omega&0&0\\ 0&0&\omega^{2}&0\\ 0&0&0&\omega^{3}\end{pmatrix}$, (17)
respectively, where $\omega=e^{2\pi i/4}=i$ and $(V_{j})^{4}=(U_{j})^{4}=\mathbbm{1}$. For the $\mathbb{Z}_{4}$ clock model, the many-body projector can be constructed in close analogy to the $\mathbb{Z}_{3}$ case, and is explicitly expressed as
$P_{q}=\frac{1}{4}\left[\mathbbm{1}+\mathbb{T}+\mathbb{T}^{2}+\mathbb{T}^{3}\right].$ (18)
Here, $\mathbb{T}=\omega^{-q}\prod_{j}U_{j}$, $q=0$ is the neutral sector, and $q=1,2,3$ are the charged sectors.
Figure 5: Fidelity zeros in the $\mathbb{Z}_{4}$ clock model. (a) Distribution of fidelity zeros in the complex-$h$ plane for $L=8$. The scatter points mark the complex field values at which the fidelity vanishes, with the critical point located around $Re(h)\approx 1$. (b) Finite-size scaling of the fidelity zeros as a function of system size $L$ from 7 to 12. The purple dots denote the complex field values with the largest real part of the fidelity zeros; the black dashed line shows the fitted curve, and the red star on the real axis marks the critical value $h_{c}=1.023$.
For the $\mathbb{Z}_{4}$ clock model, the phase transition occurs at $\mathrm{Re}(h)=1$ with critical exponent $\nu=1$ under a real field [38]. In the presence of a complex field, non-Hermitian symmetry breaking lifts the ground-state degeneracy. Specifically, the complex field splits the spectrum among the neutral sector ($q=0$), the charged sector ($q=2$), and the pair of degenerate charged sectors ($q=1,3$). As a consequence, the ground state is effectively governed by the competition between the $q=0$ and $q=2$ sectors, giving rise to fidelity zeros that provide a clear signature of the phase transition.
We numerically calculate the fidelity zeros in the complex-field plane for a system of size $L=8$, as shown in Fig. 5(a). The zeros are prominently distributed around $\mathrm{Re}(h)=1$. For $L=8$, the finite-size transition point is found at $\mathrm{Re}(h_{L})=1.0579$. Remarkably, fidelity zeros occur only in the region $\mathrm{Re}(h)<\mathrm{Re}(h_{L})$, whereas they are absent for $\mathrm{Re}(h)>\mathrm{Re}(h_{L})$. Applying the finite-size scaling analysis of Eq. (13), we determine the critical point $h_{c}=1.023$ with the correlation-length critical exponent $\nu=1$, as illustrated in Fig. 5(b). These findings provide further evidence that the fidelity-zero framework can be systematically extended to models with higher discrete symmetries.